OpenStudy (anonymous):

Calculate the enthalpy change, ΔH, for the process in which 14.5g of water is converted from liquid at 10.5∘C to vapor at 25.0∘C . For water, ΔHvap = 44.0kJ/mol at 25.0∘C and Cs = 4.18 J/(g⋅∘C) for H2O(l).

OpenStudy (anonymous):

aaronq you're my hero.

OpenStudy (aaronq):

so were assuming that water can vaporize at 25 celsius because the pressure is lower? lol I think that would change the values you were given. Do you have to account for that? If not you can just do: \(q=m*C_p*\Delta T+n*\Delta H_{vap}\) n=moles, btw; you have to use it because the value you were given is in moles.

OpenStudy (anonymous):

okay that makes sense because the little sidebar of information includes that formula.

OpenStudy (anonymous):

would it be 1 mole?

OpenStudy (aaronq):

nope, the molar mass of water is 18g/mol so just a little less, the formula for moles is: \(n=\dfrac{m}{M}\)

OpenStudy (anonymous):

okay so now i'm confused. What I did was 14.5g * 4.18 J/g-C * 14.5 + n *44.0kj/mol

OpenStudy (anonymous):

So for n it would be 18g/1 mol H2O

OpenStudy (aaronq):

\(n=\dfrac{m}{M}=\dfrac{14.5~g}{18~g/mol}\)

OpenStudy (anonymous):

oh okay that makes sense.

OpenStudy (anonymous):

i got 914.285

OpenStudy (anonymous):

but i have a feeling that i'm wrong..

OpenStudy (anonymous):

well the answer was 35.4kJ i had to figure out q for the energy of warming it up and then q for the second half and add the two together.

OpenStudy (aaronq):

you didn't take the units into account, the first part is in Joules (look at the units on the specific heat capacity), the second in kJ. \(1=(14.5* 4.18 * 14.5)J*\dfrac{1kJ}{1000~J} +(\dfrac{14.5}{18} *44.0)kJ\)

OpenStudy (aaronq):

sorry that 1 is supposed to be a q

OpenStudy (anonymous):

Yeah i realize that now =[

OpenStudy (aaronq):

damn, sorry!

OpenStudy (anonymous):

it's okay! Thank you for all your help though!